# Accelerating Graph Neural Networks on Real Processing-In-Memory Systems

### v1

@@ IV-F GNN Aggregation Performance
We draw three findings. First, SP1 and SP2 achieve small speedups over CPU (on average 1.14$\times$), since they are optimized SpMV implementations. Instead, aggregation by its nature performs SpMM, thus PyGim schemes provide significant performance speedups, on average 3.07$\times$ and up to 4.00$\times$ over CPU. Second, if the hidden size is larger, PyGim provides higher speedups than that of small hidden size. For instance, PyGim outperforms CPU by 2.37$\times$ and 3.45$\times$ for 128 and 256 hidden size, respectively. When hidden size is larger, the execution results in larger amount of data movement between memory and processors (exhibits lower arithmetic intensity), thus PIM systems can provide larger benefits thanks to their immense memory bandwidth available. Third, the best-performing PyGim scheme varies across datasets due to different connection characteristics between vertices. With 256 hidden size, the best-performing PyGim execution has 1 (RE1 and CP1) and 2 sparse partitions (RE2 and CP2) in ogbn-proteins and Reddit, respectively. This finding strengthens Recommendation 5: selection methods are needed to tune the parallelization scheme based on the characteristics of each particular input given. In this work, we leave for future work the exploration of tuning methods for the sparse/dense partitions in PyGim’s kernels. Overall, we conclude that PyGim provides significant performance benefits in aggregation for various widely-used graph datasets and hidden sizes in GNN inference.
Recommendation 8: Engineers can allocate multiple PIM-enabled memory modules based on the workload’s memory bandwidth needs, e.g., when using a larger hidden size, GNN aggregation has larger memory bandwidth needs, thus a larger number of PIM memory modules can be used.
@@ IV-G End-to-End GNN Performance
PyGim RE and CP schemes provide significant performance speedups over both their CPU counterpart and prior state-of-the-art PIM schemes (SP1 and SP2) by 3.04$\times$ (up to 3.44$\times$) and 2.46$\times$ (up to 2.68$\times$), respectively. PIM GNN execution achieves low performance with fp32 values, since their arithmetic operations are software emulated in the UPMEM PIM hardware. However, ML-oriented PIM systems are expected to be in the market soon, and will hopefully have native support for high precision data types. Instead, Fig. 15 evaluates int16 values (arithmetic operations are natively supported by PIM hardware) and shows that PyGim significantly outperforms its CPU counterpart by on average 4.23$\times$ and up to 4.94$\times$. When arithmetic operations are natively supported by the PIM hardware, as it happens for int16 data type in UPMEM PIM system, PIM GNN execution provides significant performance benefits by exploiting the immense memory bandwidth available in PIM systems. Finally, we find that when using fp32 values, PIM GNN execution achieves low performance, being worse than that of CPU by on average 39.1%. This is because UPMEM PIM hardware does not support floating-point operations, which are software emulated, thus they incur high performance overheads. However, we expect that ML-oriented PIM systems will be available in the market the next few years, and will hopefully support in hardware high precision data types.
@@ IV-H Comparison with GPU Systems
Intel Xeon 4215 7.30s 7.10s 8.04 8.26 0.63% 0.32%
UPMEM PIM 1.87s 12.74s 31.3 4.61 27.01% 18.54%
RTX3090 0.37s 0.11s 158.33 553.94 0.44% 1.56%
TABLE IV: Aggregation performance achieved in different systems.
GPU performs best in absolute execution time, since it has high compute throughput, 154$\times$ and 1431$\times$ higher than that of PIM system for int32 and fp32, respectively (Table III). GPUs were commercialized a few decades ago, and hardware architects have spent large budgets to optimize them. Instead, UPMEM PIM was released $\sim$1 year ago and is still maturing: cores are expected to run at a higher frequency in the near future (500MHz instead of 350MHz) [12]. However, PyGim aggregation achieves the largest resource utilization, being 61.3$\times$ and 11.8$\times$ larger than that of GPU for int32 and fp32, respectively. GPU execution is bottlenecked by memory bandwidth, while PyGim’s software kernels highly exploit the compute capabilities of the underlying PIM hardware. Large resource utilization is highly desirable, since it improves the cost of ownership. Thus, we conclude that PIM is a very cost-effective solution.

### v2

@@ IV-F GNN Aggregation Performance
We draw two findings. First, SP1 and SP2 achieve small speedups over CPU (on average 1.14$\times$), since they are optimized SpMV implementations. Instead, aggregation by its nature performs SpMM, thus PyGim schemes provide significant performance speedups, on average 3.07$\times$ and up to 4.00$\times$ over CPU. Second, the best-performing PyGim scheme varies across datasets due to different connection characteristics between vertices. With 256 hidden size, the best-performing PyGim execution has 1 (RE1 and CP1) and 2 sparse partitions (RE2 and CP2) in ogbn-proteins and Reddit, respectively. This finding strengthens Recommendation 5: selection methods are needed to tune the parallelization scheme based on the characteristics of each particular input given. In this work, we leave for future work the exploration of tuning methods for the sparse/dense partitions in PyGim’s kernels. Overall, we conclude that PyGim provides significant performance benefits in aggregation for various widely-used graph datasets and hidden sizes in GNN inference.
@@ IV-G End-to-End GNN Performance
PyGim RE and CP schemes provide significant performance speedups over both their CPU counterpart and prior state-of-the-art PIM schemes (SP1 and SP2) by 3.04$\times$ (up to 3.44$\times$) and 2.46$\times$ (up to 2.68$\times$), respectively. PIM GNN execution achieves low performance with fp32 values, since their arithmetic operations are software emulated in the UPMEM PIM hardware. However, ML-oriented PIM systems are expected to be in the market soon, and will hopefully have native support for high precision data types. Instead, Fig. 15 evaluates int16 values (arithmetic operations are natively supported by PIM hardware) and shows that PyGim significantly outperforms its CPU counterpart by on average 4.03$\times$ and up to 4.63$\times$. When arithmetic operations are natively supported by the PIM hardware, as it happens for int16 data type in UPMEM PIM system, PIM GNN execution provides significant performance benefits by exploiting the immense memory bandwidth available in PIM systems. Finally, we find that when using fp32 values, PIM GNN execution achieves low performance, being worse than that of CPU by on average 41.6%. This is because UPMEM PIM hardware does not support floating-point operations, which are software emulated, thus they incur high performance overheads. However, we expect that ML-oriented PIM systems will be available in the market the next few years, and will hopefully support in hardware high precision data types.
@@ IV-H Comparison with GPU Systems
Intel Xeon 4215 6.93 11.77 4.23 2.49 0.66% 0.19%
UPMEM PIM 2.03 13.02 14.43 2.25 12.45% 9.07%
GPU RTX3090 0.11 0.11 279.12 275.86 1.56% 0.77%
TABLE IV: Aggregation performance for Reddit in different systems.
GPU performs best in absolute execution time, since it has high compute throughput, 154$\times$ and 1431$\times$ higher than that of PIM system for int32 and fp32, respectively (Table III). GPUs were commercialized a few decades ago, and hardware architects have spent large budgets to optimize them. Instead, UPMEM PIM was released $\sim$1 year ago and is still maturing: cores are expected to run at a higher frequency in the near future (500MHz instead of 350MHz) [12]. However, PyGim aggregation achieves the largest resource utilization, being 7.9$\times$ and 11.7$\times$ larger than that of GPU for int32 and fp32, respectively. GPU execution is bottlenecked by memory bandwidth, while PyGim’s software kernels highly exploit the compute capabilities of the underlying PIM hardware. Large resource utilization is highly desirable, since it improves the cost of ownership. Thus, we conclude that PIM is a very cost-effective solution.